# MegaScale-MoE: Large-Scale Communication-Efficient Training of Mixture-of-Experts Models in Production

### v2

@@ 4.2 Intra-operator Overlap
A widely adopted solution [17, 48, 50] is to decompose operators into smaller parallel ones to enable pipelining by executing them on separate CUDA streams. However, this approach introduces non-negligible overhead: complex stream control, involving host interference and causing random bubbles due to the non-deterministic feature of CPU control; imperfect tail computation, increasing overall computation latency.
To address the above issues, we use intra-operator overlap to exploit the parallelism for communication and computation operators with direct dependency. The core idea is to fuse these operators and break down the workloads into tiles. Following prior work [15, 5, 51, 53], we implement barriers in device memory between communication and computation operators. These barriers allow fine-grained notifications between computation and communication at the tile level. Additionally, as the barriers reside in device memory, they eliminate the need for host interference, enabling further performance improvements. We implement two types of kernels, overlapping with GEMMs and overlapping with MoE GroupedGEMMs, for the attention and FFN modules, respectively.
@@
For A2A+GEMM and GEMM+A2A, we allocate a small number of SMs for communication as all-to-all is more complex than all-gather and reduce-scatter. The number of SMs for communication is tuned to make communication and computation exhibit similar latency. Moreover, multiple ranks may simultaneously read from or write to the same device, potentially causing contention in NVLink. To mitigate this, we apply swizzling [5, 51, 53] to reorder tile communication and computation so that the arrival of communication tiles aligns with the pace of computation tiles.
@@ 5 Communication Compression
6 Evaluation
| System | #GPUs | Iteration Time (s) | Throughput (tokens/s) | Training Time for 1T Tokens (days) |
| Megatron-LM | 240 | 39.94 | 151.1k | 76.61 |
| 480 | 19.56 | 301.1k | 38.38 | |
| 720 | 13.70 | 430.5k | 26.88 | |
| 960 | 10.82 | 550.2k | 21.23 | |
| 1440 | 7.90 | 746.6k | 15.50 | |
| MegaScale-MoE | 240 | 21.61 | 272.9k (1.81) | 42.41 |
| 480 | 11.83 | 498.6k (1.65) | 23.21 | |
| 720 | 7.97 | 740.1k (1.72) | 15.64 | |
| 960 | 6.12 | 963.8k (1.77) | 12.01 | |
@@ 6 Evaluation
6.1 Training Performance
MegaScale-MoE is built on top of Megatron-LM [46], a state-of-the-art open-source LLM training system that supports 3D parallelism strategies and is continuously updated to incorporate the latest optimizations from the community. Our evaluation uses the Megatron-LM on GitHub [30] with commit hash f1f03922, selected for its stability at the commencement of our experiments months ago. For fair comparison, we use the same global batch size for Megatron-LM and MegaScale-MoE and choose the optimal parallelism configurations for the two systems, respectively. Specifically, MegaScale-MoE employs SP attention and EP within each node, while Megatron-LM adopts TP within each node, with both systems configured with a PP size of 15. Sequence length is 8,192 and vocabulary size is 65,536.
@@ 6.1 Training Performance
Figure 11 presents the weak-scaling training performance of Megatron-LM and MegaScale-MoE on the same model. We scale the global batch size from 360 to 1,080 in proportion to the number of GPUs (from 480 to 1,440). MegaScale-MoE achieves a 1.74-1.79 training throughput compared to Megatron-LM. As the scale increases, Megatron-LM’s throughput drops by 2.74% due to increased communication overhead, whereas MegaScale-MoE exhibits near-linear scalability, benefiting from comprehensive communication overlap.
| GPU | Compute Cap- ability (TFLOPS) | Memory Spec. | NVLink Bw. (GB/s) | |
| Cap. (GB) | Bw. (TB/s) | |||
| H800 | 989 | 80 | 3.4 | 400 |
| A100 | 312 | 80 | 2.0 | 600 |
| H20 | 148 | 96 | 4.0 | 900 |
@@
6.2 Ablation Study
Parallelism strategy. We first compare the training efficiency under various intra-node parallelism strategies using a single node with eight NVIDIA H800-SXM GPUs. We denote parallelism strategies as X+Y, where X represents the parallelism strategy for attention, and Y corresponds to that for experts. The available parallelism strategies for attention include TP and our SP, whereas for experts, the choices are TP and EP. To isolate the performance benefits of optimized parallelism, we disable other system optimizations.
@@ 7 Experience
Deployment experience. MegaScale-MoE has been deployed in our production environment and is responsible for the majority of large-scale MoE training tasks within our company. It enables the training of models with trillions of parameters, supports single training jobs scaling beyond 10,000 GPUs, with individual training tasks running for several months. By combining the aforementioned techniques, MegaScale-MoE minimizes idle communication time and optimizes memory usage in MoE training without compromising model performance, ultimately saving millions of GPU hours in large-scale MoE training. Figure 19 shows the model convergence from a real production run, which trains a proprietary MoE model with 200B parameters, 20B activated for each token. This run uses over 10,000 GPUs and lasts for months. The loss continues to converge with a stable training process.
@@
Beyond ensuring training convergence, we introduce additional engineering optimizations. Existing FP8 training implementations [48, 23] store model parameters in BF16, requiring frequent FP8 conversion for GEMM computations, adding casting and transpose overhead. To address this, we use a multi-precision optimizer to store model parameters directly in FP8, while keeping main parameters in FP32 with separate buffers for different data types. This lowers memory consumption and halves parameter all-gather communication in data parallelism.
@@
| comm_time | (5) | |||
| comp_time | (6) |
@@
8 Related Work
@@
Large model training. LLM research has led to the development of scalable, efficient, and robust training techniques to meet the substantial computational demands of these models. DeepSpeed [41] features the Zero Redundancy Optimizer (ZeRO) [38, 39, 42], which shards model parameters, gradients, and optimizer states across participating GPUs in data parallelism, enabling the scaling of LLMs with manageable memory consumption. Megatron-LM [46] focuses on intra-layer model parallelism techniques, partitioning the parameters and computation of each layer. Pipeline parallelism assigns the parameters and computation of a contiguous subset of layers to each GPU[12, 31], breaks a batch into micro-batches, and processes the micro-batches in a pipelined fashion. MegaScale [17] shows how combining tensor, pipeline, and data parallelism can be an efficient strategy to train large multi-billion parameter models at unprecedented scale.
@@
Mixture-of-Expert training. To address the computational challenges of training advanced neural networks, the machine learning field has increasingly adopted Mixture-of-Experts architectures. Subsequently, a number of parallel deep learning frameworks have been proposed for training or running inference on MoEs on multi-GPU clusters. DeepSpeed-MoE [40] significantly reduces training costs through model architecture designs and compression techniques. HetuMoE [34] utilizes a hierarchical all-to-all communication strategy to achieve performance speedup. SE-MoE [45] distinguishes itself by focusing on scalable and efficient training with heterogeneous resources like CPU memory and SSDs. Furthermore, Tutel [13] offers a dynamic solution for MoE models, employing adaptive parallelism and pipelining. FasterMoE [11] introduces a comprehensive suite of optimizations such as dynamic shadowing, fine-grained scheduling, and congestion-avoiding expert selection strategies. Janus [28] proposes a data-centric paradigm shift for MoE models, aiming to lower communication demands and boost training efficiency. Recently, DeepSeek-V3 [25] proposes a series of algorithm-system co-optimizations, such as efficient cross-node all-to-all and DualPipe, to enhance the efficiency of large-scale MoE training. We are also exploring algorithm-system co-design to further reduce training costs.
@@
Long-context training. While Megatron-LM [46, 18] opts to partition only specific operations along the sequence dimension, various methods of sequence parallelism [19, 27, 22] have been explored for training models requiring long contexts. The Blockwise Parallel Transformer [26] method implements blockwise computation of self-attention and the fusion of FFNs based on online softmax calculations. Ring Attention [27, 22] introduces a ring-style communication mechanism integrated with self-attention calculations, facilitating the exchange of key and value chunks. We adopt the all-to-all style of SP attention from DeepSpeed Ulysses [14], which partitions attention by heads rather than sequence length, due to its reduced communication volume and balanced computation pattern.
@@
Communication-computation overlap. Several frameworks [10, 21, 29, 37, 52] focus on overlapping communication with computation in distributed deep learning training with a single parallelism strategy. Some compiler-style work [15, 50, 35] provides fine-grained overlap among kernels, but excessive partitioning of GEMM kernels can result in low GPU utilization. Centauri [6] enhances communication overlap for LLM training with 3D parallelism by communication partitioning and hierarchical scheduling. Similar to Centauri, our inter-operator communication overlap hides communication within independent computation by reordering operators. We further conceal communication on critical paths through intra-operator overlap, without compromising GPU utilization. DualPipe, proposed by DeepSeek-V3 [25], overlaps communication and computation within different forward and backward chunks and requires storing 2 the model parameters. In contrast, MegaScale-MoE achieves this overlap within a single forward or backward chunk without incurring additional memory overhead.
9 Conclusion
@@ Appendix A Appendix
A.1 Hierarchical Communication for Parameter Synchronization
@@
A.2 Selective Activation Rematerialization
which we have reduced to

### v3

@@ 4.2. Intra-operator Overlap
A widely adopted solution (Jiang et al., 2024a; TransformerEngine, 2025; Wang et al., 2022) is to decompose operators into smaller parallel ones to enable pipelining by executing them on separate CUDA streams. However, this approach introduces non-negligible overhead: complex stream control, involving host interference and causing random bubbles due to the non-deterministic feature of CPU control; imperfect tail computation, increasing overall computation latency.
To address the above issues, we adopt intra-operator overlap to parallelize communication and computation operators with direct dependencies. The core idea is to fuse these operators and break down the workloads into tiles. Following prior work (Jangda et al., 2022; Chang et al., 2024; Zhang et al., 2025a; Zheng et al., 2025), we implement barriers in device memory between communication and computation operators. These barriers enable fine-grained tile-level notifications and remove the need for host interference, further improving training performance. We implement two types of kernels, overlapping with GEMMs and overlapping with MoE GroupedGEMMs, for the attention and FFN modules, respectively.
@@
For A2A+GEMM and GEMM+A2A, we allocate a small number of SMs for communication as all-to-all is more complex than all-gather and reduce-scatter. The number of SMs for communication is tuned to make communication and computation exhibit similar latency. Moreover, multiple ranks may simultaneously read from or write to the same device, potentially causing contention in NVLink. To mitigate this, we apply swizzling (Chang et al., 2024; Zhang et al., 2025a; Zheng et al., 2025) to reorder tile communication and computation so that the arrival of communication tiles aligns with the pace of computation tiles.
@@ 5. Communication Compression
6. Evaluation
| Megatron-LM | 240 | 39.94 | 151.1k | 76.61 | ||||||||
| 480 | 19.56 | 301.1k | 38.38 | |||||||||
| 720 | 13.70 | 430.5k | 26.88 | |||||||||
| 960 | 10.82 | 550.2k | 21.23 | |||||||||
| 1440 | 7.90 | 746.6k | 15.50 | |||||||||
| MegaScale-MoE | 240 | 21.61 | 272.9k (1.81) | 42.41 | ||||||||
| 480 | 11.83 | 498.6k (1.65) | 23.21 | |||||||||
| 720 | 7.97 | 740.1k (1.72) | 15.64 | |||||||||
| 960 | 6.12 | 963.8k (1.77) | 12.01 | |||||||||
@@ 6. Evaluation
6.1. Training Performance
MegaScale-MoE is built on top of Megatron-LM (Shoeybi et al., 2019), a state-of-the-art open-source LLM training system that supports 3D parallelism strategies and is continuously updated to incorporate the latest optimizations from the community. Our evaluation uses the Megatron-LM on GitHub (Megatron-LM, 2025) with commit hash f1f03922, selected for its stability at the commencement of our experiments months ago. For fair comparison, we use the same global batch size for Megatron-LM and MegaScale-MoE and choose the optimal parallelism configurations for the two systems, respectively. Specifically, MegaScale-MoE employs SP attention and EP within each node, while Megatron-LM adopts TP within each node, with both systems configured with a PP size of 15. We tune the configuration of Megatron-LM to meet its requirement of a uniform TP size across all components. As discussed in §3.1, for Megatron-LM, a TP size of 1 leads to a prohibitive 8 activation memory (addressable only with slow recomputation via gradient checkpointing), while a TP size of 8 forces EP to operate across nodes, incurring more communication costs than PP. Notably, both systems in the evaluation enable the communication-computation overlap techniques from MegaScale (Jiang et al., 2024a) for data and pipeline parallelism. Therefore, the communication overhead mainly comes from intra-node model parallelism, e.g. TP, SP and EP. Sequence length is 8,192 and vocabulary size is 65,536.
@@ 6.1. Training Performance
Figure 12 presents the weak-scaling training performance of Megatron-LM and MegaScale-MoE on the same model. We scale the global batch size from 360 to 1,080 in proportion to the number of GPUs (from 480 to 1,440). MegaScale-MoE achieves a 1.74-1.79 training throughput compared to Megatron-LM. As the scale increases, Megatron-LM’s throughput degrades by 2.74% due to increased communication overhead. In contrast, MegaScale-MoE exhibits near-linear scalability, with its throughput declining by only 0.2%, benefiting from comprehensive communication-computation overlap.
| GPU |
Memory Spec. |
| Cap. (GB) | Bw. (TB/s) | |||||||
| H800 | 989 | 80 | 3.4 | 400 | ||||
| A100 | 312 | 80 | 2.0 | 600 | ||||
| H20 | 148 | 96 | 4.0 | 900 | ||||
@@
6.2. Ablation Study
| 1 | baseline | 1 | ||||||
| 2 | (1) with SP+EP | 1.13 | +13% | |||||
| 3 | (2) with inter-operator overlap | 1.22 | +9% | |||||
| 4 | (3) with intra-operator overlap | 1.28 | +6% |
We evaluate the effectiveness of the optimization techniques of MegaScale-MoE. First, we conduct an experiment about systematic breakdown by incrementally enabling each technique to isolate its contribution to the overall performance. Table 5 shows the throughput improvement breakdown with different optimizations when training the 352B MoE model on 240 GPUs with a global batch size of 720. The baseline is a version of MegaScale-MoE that adopts TP for both attention and FFNs and disables communication-computation overlap. First, by applying communication-efficient strategies—namely, SP for attention and EP for experts—we achieve an initial 13% throughput improvement over this baseline. We then target the primary bottleneck in large-scale MoE training: communication overhead. Our inter-operator and intra-operator overlap methods effectively hide these costs, further accelerating training by an additional 9% and 6%, respectively.
Following the systematic breakdown, we perform ablation studies on each component, varying a single setting at a time while keeping all others constant, to gain deeper insights into its behavior.
Parallelism strategy. We compare the training efficiency under various intra-node parallelism strategies using a single node with eight NVIDIA H800-SXM GPUs. We denote parallelism strategies as X+Y, where X represents the parallelism strategy for attention, and Y corresponds to that for experts. The available parallelism strategies for attention include TP and our SP, whereas for experts, the choices are TP and EP. To isolate the performance benefits of optimized parallelism, we disable other system optimizations.
@@ 7. Experience
Deployment experience. MegaScale-MoE has been deployed in our production environment and is responsible for the majority of large-scale MoE training tasks within our company. It enables the training of models with trillions of parameters, supports single training jobs scaling beyond 10,000 GPUs, with individual training tasks running for several months. By combining the aforementioned techniques, MegaScale-MoE minimizes idle communication time and optimizes memory usage in MoE training without compromising model performance, ultimately saving millions of GPU hours in large-scale MoE training. Figure 20 shows the model convergence from a real production job, which trains a proprietary MoE model with 200B parameters, 20B activated for each token. This job uses over 10,000 GPUs and lasts for months. The loss continues to converge with a stable training process.
@@
Beyond ensuring training convergence, we introduce additional engineering optimizations. Existing FP8 training implementations (TransformerEngine, 2025; Liang et al., 2024) store model parameters in BF16, requiring frequent FP8 conversion for GEMM computations, adding casting and transpose overhead. To address this, we use a multi-precision optimizer to store model parameters directly in FP8, while keeping main parameters in FP32 with separate buffers for different data types. This lowers memory consumption and halves parameter all-gather communication in data parallelism.
@@
| (5) | comm_time | |||
| (6) | comp_time |
@@
MoE vs. dense model training. In our continued efforts to optimize MoE model training, we have identified several critical distinctions from the training of dense models. In a dense Transformer layer, optimization efforts are concentrated on self-attention and GEMMs. The former is often accelerated by techniques like FlashAttention (Dao et al., 2022), while the latter, as a dense computation, generally achieves high utilization on the GPU’s parallel processing units. In contrast, as shown in Figure 13a, the combined runtime of attention and GroupedGEMM accounts for only about one-third of a layer’s execution time. The remainder is consumed by communication and other operators. While MegaScale-MoE effectively addresses the communication overhead, we observe that the computational operators in MoE models, which are inherently more complex than their dense counterparts, also introduce performance degradation. Specifically, they are a primary source of stragglers for three main reasons:
First, the intermediate dimension of each expert is smaller than the FFN layer in a dense model. To efficiently process computations for multiple experts concurrently, GroupedGEMM employs a single CUDA kernel for numerous small matrix multiplications. The resource usage of this kernel—including shared memory, L1 cache, and number of threads—is finely controlled via cuFuncSetAttribute. This granular control, however, can introduce synchronization delays. Second, due to the imbalanced number of tokens routed to each expert, the inputs and outputs for GroupedGEMM are dynamically shaped tensors. The frequent allocation and deallocation of these tensors exacerbate GPU memory fragmentation. Third, the MoE gating mechanism involves a multitude of small operators for tasks like calculating routing scores and communicating routing decisions. Jitter in CPU performance can delay the launch of these kernels to the point where the launch latency exceeds their actual execution time on the GPU, creating pipeline bubbles.
8. Related Work
@@
Large model training. LLM research has led to the development of scalable, efficient, and robust training techniques (Rasley et al., 2020; Shoeybi et al., 2019; Jiang et al., 2024a; Zhang et al., 2025b) to meet the substantial computational demands of these models. DeepSpeed (Rasley et al., 2020) features the Zero Redundancy Optimizer (ZeRO) (Rajbhandari et al., 2020, 2021; Ren et al., 2021), which shards model parameters, gradients, and optimizer states across participating GPUs in data parallelism, enabling the scaling of LLMs with manageable memory consumption. Megatron-LM (Shoeybi et al., 2019) focuses on intra-layer model parallelism techniques, partitioning the parameters and computation of each layer. Pipeline parallelism assigns the parameters and computation of a contiguous subset of layers to each GPU(Huang et al., 2019; Narayanan et al., 2019), breaks a batch into micro-batches, and processes the micro-batches in a pipelined fashion. MegaScale (Jiang et al., 2024a) shows how combining tensor, pipeline, and data parallelism can be an efficient strategy to train large multi-billion parameter models at unprecedented scale.
@@
Mixture-of-Expert training. To address the computational challenges of training advanced neural networks, the machine learning field has increasingly adopted Mixture-of-Experts architectures. Subsequently, a number of deep learning frameworks have been proposed for training or running inference on MoEs on multi-GPU clusters. DeepSpeed-MoE (Rajbhandari et al., 2022) significantly reduces training costs through model architecture designs and compression techniques. HetuMoE (Nie et al., 2022) utilizes a hierarchical all-to-all communication strategy to achieve performance speedup. SE-MoE (Shen et al., 2022) distinguishes itself by focusing on scalable and efficient training with heterogeneous resources like CPU memory and SSDs. FasterMoE (He et al., 2022) introduces a comprehensive suite of optimizations such as dynamic shadowing, fine-grained scheduling, and congestion-avoiding expert selection strategies. Janus (Liu et al., 2023a) proposes a data-centric paradigm shift for MoE models, aiming to lower communication demands and boost training efficiency. Tutel (Hwang et al., 2023) offers a dynamic solution for MoE models, employing adaptive parallelism and pipelining. However, its dynamic parallelism switching and hierarchical all-to-all can cause significant overheads for models with hundreds of billions of parameters. To avoid such overhead, latest MoE training systems (Liu et al., 2024a, b) use auxiliary loss or routing bias for load balancing and limit cross-node token dispatch. By mapping each MoE layer to intra-node, MegaScale-MoE eliminates cross-node token dispatch.
Recently, DeepSeek-V3 (Liu et al., 2024b) introduced two key optimizations for training production-scale MoE models: DeepEP, for high-performance cross-node all-to-all communication, and DualPipe, for overlapping communication with computation. Due to the relatively low cross-node InfiniBand bandwidth, DeepEP limits the token dispatch to a maximum of 4 nodes to maintain a constant cross-node communication volume, restricting its routing flexibility. In contrast, MegaScale-MoE places each MoE layer intra-node to ensure efficient routing to any top-k experts. DualPipe leverages pipeline parallelism for communication-computation overlap across different micro-batches, which requires storing 2 the model parameters. In contrast, MegaScale-MoE’s overlap occurs within a single micro-batch’s forward or backward pass, incurring no additional memory overhead and remaining compatible with systems both with and without pipeline parallelism.
@@
Long-context training. While Megatron-LM (Shoeybi et al., 2019; Korthikanti et al., 2023) opts to partition only specific operations along the sequence dimension, various methods of sequence parallelism (Li et al., 2024; Liu et al., 2023b; Li et al., 2021; Gu et al., 2024) have been explored for training models requiring long contexts. The Blockwise Parallel Transformer (Liu and Abbeel, 2024) method implements blockwise computation of self-attention and the fusion of FFNs based on online softmax calculations. Ring Attention (Liu et al., 2023b; Li et al., 2021) introduces a ring-style communication mechanism integrated with self-attention calculations, facilitating the exchange of key and value chunks. We adopt the all-to-all style of SP attention from DeepSpeed Ulysses (Jacobs et al., 2023), which partitions attention by heads rather than sequence length, due to its reduced communication volume and balanced computation pattern.
@@
Communication-computation overlap. Several frameworks (Hashemi et al., 2019; Li et al., 2020; Mahajan et al., 2023; Peng et al., 2019; Zhao et al., 2023) focus on overlapping communication with computation in distributed deep learning training with a single parallelism strategy. Some compiler-style work (Jangda et al., 2022; Wang et al., 2022; Pati et al., 2024) provides fine-grained overlap among kernels, but excessive partitioning of GEMM kernels can result in low GPU utilization. Centauri (Chen et al., 2024) enhances communication overlap for LLM training with 3D parallelism by communication partitioning and hierarchical scheduling. Similar to Centauri, our inter-operator communication overlap hides communication within independent computation by reordering operators. We further conceal communication on critical paths through intra-operator overlap, without compromising GPU utilization.
9. Conclusion
@@
Acknowledgements. We thank our shepherd, Cheng Li, and the anonymous reviewers for their valuable feedback and suggestions. This work was supported in part by the National Key Research and Development Program of China under Grant 2022YFB4500700, the Scientific Research Innovation Capability Support Project for Young Faculty under Grant ZYGXQNJSKYCXNLZCXM-I1, the Fundamental Research Funds for the Central Universities, Peking University, and the National Natural Science Foundation of China under Grant 62172008 and Grant 62325201. Xin Jin and Xin Liu are the corresponding authors. Chao Jin, Xuanzhe Liu, and Xin Jin are also with the Key Laboratory of High Confidence Software Technologies (Peking University), Ministry of Education.
@@ Appendix A Appendix
A.1. Hierarchical Communication for Parameter Synchronization